# Universal Exotic Dynamics in Critical Mesoscopic Systems: Simulating the Square Root of Avogadro’s Number of Spins

### v1

@@ Appendix C Correlation function analysis
The coherence length was calculated from $\xi(t)=I_{2}(t)/I_{1}(t)$. The large system sizes we use enabled the high accuracy of $C(r;t)$, allowing us to reach $\xi$ values up to $1133.3(4)$. Specifically, the integration of Eq. (13) involved three major steps: (i) A cut-off distance, $r_{\rm cut}$, is determined such that $C(r_{\rm cut}+1;t)$ is smaller than three times its error, thereby establishing an upper integration cutoff based on the accuracy of our data. (ii) Next, the maximum of $r^{2}C(r;t)$, located at $r^{\ast}$, is found. From its value, ${r^{\ast}}^{2}C(r^{\ast};t)$, the characteristic distances $r_{\rm min}$ and $r_{\rm max}$ are defined as the first $r$ where $r_{\rm min}^{2}C(r_{\rm min};t)<0.9{r^{\ast}}^{2}C(r^{\ast};t)$ and $r_{\rm max}^{2}C(r_{\rm max};t)<{r^{\ast}}^{2}C(r^{\ast};t)/3$, respectively. We require that $r^{\ast}<r_{\rm min}<r_{\rm max}<r_{\rm cut}$; if this condition is not met, we increase our statistics by performing additional simulations. These characteristic distances are then used to calculate the integrals of Eq. (13). (iii) If $r_{\rm max}-r_{\rm min}\leq 8$, a numerical integration of $C(r;t)$ is performed up to $r_{\rm cut}$. However, if $r_{\rm max}-r_{\rm min}>8$, $C(r;t)$ is fitted in the interval $[r_{\rm min},r_{\rm max}]$ to the function
@@ Appendix D Fitting procedures
Given our data, the exponent $z$ can be calculated using the two methods discussed in the main text, which are reiterated here for the reader’s convenience. Starting from the equation $\frac{d\log{(t)}}{d\log{(\xi)}}=z+a/\xi^{\omega}$ for a finite system, and integrating, we obtain the following relation for the coherence length
@@ Appendix E Capturing the thermodynamic limit
In the out-of-equilibrium regime, correlations at a distance $r$ decay exponentially as $[r/\xi(t)]^{a}$, where the exponent $a>1$. This suggests that finite-size effects should decay exponentially as $[L/\xi(t)]^{a}$. Consequently, significant finite-size effects in out-of-equilibrium dynamics are expected only when the coherence length $\xi(t)$ becomes larger than a fraction of the system size $L$. A typical threshold is $1/7$, meaning that finite-size effects are often considered negligible when $\xi(t)<L/7$ Belletti et al. (2008)). However, this rule of thumb overlooks a more important consideration: namely, are the finite-size effects in our simulations significantly smaller than the statistical errors? In other words, the acceptable size of finite-size effects is not determined a priori but is instead set by the statistical errors in the simulation. In our case, these errors are particularly small due to the large system sizes used in our simulations ($L=2^{19}$ for the Blume-Capel model and up to $L=2^{22}$ for Ising model). Given that we limited $\xi(t)$ to be no larger than approximately $10^{3}$ lattice spacings, we are in the range where $L/\xi\approx 4\times 10^{3}$, which suggests that finite-size effects are unlikely to be significant. Nonetheless, we have performed a quantitative check, which we describe next.
@@
Figure 8 shows the difference in the values of $C(r,t)$ obtained from simulations for $L=2^{20}$ and $L=2^{22}$, for four values of the spatial separation $r$: the energy ($r=1$), $r=10$, $r=100$ and $r=1002$. The differences are consistent with zero in all four cases. Notable time correlations are visible in Fig. 8, where the difference remains either positive or negative for a significant period. However, deviations from zero at any given time are well within the statistical errors.

### v2

@@ Appendix C Correlation function analysis
The coherence length was calculated from $\xi(t)=I_{2}(t)/I_{1}(t)$. The large system sizes we use enabled the high accuracy of $C(r;t)$, allowing us to reach $\xi$ values up to 1133.3(4)1133.3(4)1133.3 ( 4 ). Specifically, the integration of Eq. (13) involved three major steps: (i) A cut-off distance, $r_{\rm cut}$, is determined such that $C(r_{\rm cut}+1;t)$ is smaller than three times its error, thereby establishing an upper integration cutoff based on the accuracy of our data. (ii) Next, the maximum of $r^{2}C(r;t)$, located at $r^{\ast}$, is found. From its value, ${r^{\ast}}^{2}C(r^{\ast};t)$, the characteristic distances $r_{\rm min}$ and $r_{\rm max}$ are defined as the first $r$ where $r_{\rm min}^{2}C(r_{\rm min};t)<0.9{r^{\ast}}^{2}C(r^{\ast};t)$ and $r_{\rm max}^{2}C(r_{\rm max};t)<{r^{\ast}}^{2}C(r^{\ast};t)/3$, respectively. We require that $r^{\ast}<r_{\rm min}<r_{\rm max}<r_{\rm cut}$; if this condition is not met, we increase our statistics by performing additional simulations. These characteristic distances are then used to calculate the integrals of Eq. (13). (iii) If $r_{\rm max}-r_{\rm min}\leq 8$, a numerical integration of $C(r;t)$ is performed up to $r_{\rm cut}$. However, if $r_{\rm max}-r_{\rm min}>8$, $C(r;t)$ is fitted in the interval $[r_{\rm min},r_{\rm max}]$ to the function
@@ Appendix D Fitting procedures
Given our data, the exponent $z$ can be calculated using the two methods discussed in the main text, which are reiterated here for the reader’s convenience. Starting from the equation $\frac{d\log{(t)}}{d\log{(\xi)}}=z+a\xi^{-\omega}$ for a finite system, and integrating, we obtain the following relation for the coherence length
@@ Appendix E Capturing the thermodynamic limit
In the out-of-equilibrium regime, correlations at a distance $r$ decay exponentially as $[r/\xi(t)]^{a}$, where the exponent $a>1$. This suggests that finite-size effects should decay exponentially as $[L/\xi(t)]^{a}$. Consequently, significant finite-size effects in out-of-equilibrium dynamics are expected only when the coherence length $\xi(t)$ becomes larger than a fraction of the system size $L$. A typical threshold is 1/71/71 / 7, meaning that finite-size effects are often considered negligible when $\xi(t)<L/7$ Belletti et al. (2008)). However, this rule of thumb overlooks a more important consideration: namely, are the finite-size effects in our simulations significantly smaller than the statistical errors? In other words, the acceptable size of finite-size effects is not determined a priori but is instead set by the statistical errors in the simulation. In our case, these errors are particularly small due to the large system sizes used in our simulations ($L=2^{19}$ for the Blume-Capel model and up to $L=2^{22}$ for Ising model). Given that we limited $\xi(t)$ to be no larger than approximately $10^{3}$ lattice spacings, we are in the range where $L/\xi\approx 4\times 10^{3}$, which suggests that finite-size effects are unlikely to be significant. Nonetheless, we have performed a quantitative check, which we describe next.
Figure 8 shows the difference in the values of $C(r,t)$ obtained from simulations for $L=2^{20}$ and $L=2^{22}$, for four values of the spatial separation $r$: the energy ($r=1$), $r=10$, $r=100$, and $r=1002$. The differences are consistent with zero in all four cases. Notable time correlations are visible in Fig. 8, where the difference remains either positive or negative for a significant period. However, deviations from zero at any given time are well within the statistical errors.